# Recognition through Reasoning: Reinforcing Image Geo-localization with Large Vision-Language Models

### v1

@@ 5 Discussion
Toward Fine-Grained Geo-localization: Limits of Pure Reasoning. While our reasoning-aware framework achieves strong performance at the country and city levels, its effectiveness diminishes when tasked with fine-grained, coordinate-level localization. This limitation originates from the inherent nature of the reasoning process: predictions are based on high-level semantic cues such as language, architectural style, or vegetation, which often lack the spatial specificity required to differentiate between closely situated locations. For example, multiple European cities may share similar visual patterns, such as Mediterranean-style architecture, the presence of European Union flags, or public signage in English, which makes it difficult for the model to resolve fine-grained geographic ambiguities through reasoning alone. In such cases, even accurate reasoning can only narrow down a broad region but cannot pinpoint an exact location. This highlights a key challenge in reasoning-centric geo-localization: the lack of precise visual-geographic anchoring. To overcome this limitation, future work may explore hybrid approaches that combine reasoning to constrain the candidate region, followed by local feature-based retrieval within that region to achieve coordinate-level precision.
Beyond Scale Alone: Data Efficiency in Reasoning-aware Training. Our experiments show that training GLOBE on just 33K high-quality, reasoning-aware samples (MP16-Reason) achieves performance comparable to, and sometimes exceeding, models trained on millions of generic image-text pairs. This highlights that for reasoning-centric tasks, targeted supervision can be more effective than sheer data scale. Our results suggest that aligning supervision with task-specific reasoning offers a more data-efficient path forward for LVLM training.
@@ 6 Conclusion
In this paper, we present a novel reasoning-driven pipeline for image geo-localization by leveraging LVLMs. To address the limitations of existing datasets and training paradigms, we introduce MP16-Reason, a high-quality dataset constructed from diverse social media images and enriched with automatically distilled locatability labels and reasoning trajectories. Building upon this dataset, we propose GLOBE, an LVLM trained via GRPO-based reinforcement learning, which jointly improves three core aspects of geo-localization: locatability assessment, visual clue reasoning, and geo-location recognition. In contrast to SFT, our GRPO-based training framework directly optimizes reasoning quality through structured reward signals, leading to substantial gains in both interpretability and localization accuracy. Empirical results show that GLOBE, using only 33K data, achieves performance comparable to or better than state-of-the-art methods trained on millions of samples.
@@ A.1.1 Prompt Template
To ensure fairness and consistency in both training and evaluation, we standardize the input format for all LVLMs. Specifically, images are fed into the model along with a unified prompt template. This prompt is designed to guide the model toward reasoning about geographic clues in a consistent manner across different stages and model variants. During both fine-tuning and testing, we use the same prompt format to minimize prompt-induced variance and allow for fair comparison of model performance.
Prompt Template You are a geolocation expert. You are participating in a geolocation challenge. Based on the provided image: 1. Carefully analyze the image for clues about its location (architecture, signage, vegetation, terrain, etc.) 2. Think step-by-step about what country, and city this is likely to be in and why Your final answer include these two lines somewhere in your response: country: [country name] city: [city name] You MUST output the thinking process in <think> </think> and give answer in <answer> </answer> tags.
@@ A.1.3 Hyper-parameter Settings
A.2 Limitation

### v2

@@ 5 Discussion
Toward Fine-Grained Geo-localization: Limits of Pure Reasoning. While our reasoning-driven framework achieves strong performance at the country and city levels, its effectiveness diminishes when tasked with fine-grained, coordinate-level localization. This limitation originates from the inherent nature of the reasoning process: predictions are based on high-level semantic cues such as language, architectural style, or vegetation, which often lack the spatial specificity required to differentiate between closely situated locations. For example, multiple European cities may share similar visual patterns, such as Mediterranean-style architecture, the presence of European Union flags, or public signage in English, which makes it difficult for the model to resolve fine-grained geographic ambiguities through reasoning alone. In such cases, even accurate reasoning can only narrow down a broad region but cannot pinpoint an exact location. This highlights a key challenge in reasoning-driven geo-localization: the lack of precise visual-geographic anchoring. To overcome this limitation, future work may explore hybrid approaches that combine reasoning to constrain the candidate region, followed by local feature-based retrieval within that region to achieve coordinate-level precision.
Beyond Scale Alone: Data Efficiency in Reasoning-driven Training. Our experiments show that training GLOBE on just 33K high-quality, reasoning-oriented samples (MP16-Reason) achieves performance comparable to, and sometimes exceeding, models trained on millions of generic image-text pairs. This highlights that for reasoning-driven tasks, targeted supervision can be more effective than sheer data scale. Our results suggest that aligning supervision with task-specific reasoning offers a more data-efficient path forward for LVLM training.
@@ 6 Conclusion
In this paper, we present a novel reasoning-driven pipeline for image geo-localization by leveraging LVLMs. To address the limitations of existing datasets and training paradigms, we introduce MP16-Reason, a high-quality dataset constructed from diverse social media images and enriched with automatically distilled localizability labels and reasoning trajectories. Building upon this dataset, we propose GLOBE, an LVLM trained via GRPO-based reinforcement learning, which jointly improves three core aspects of geo-localization: localizability assessment, visual-cue reasoning, and geo-location recognition. In contrast to SFT, our GRPO-based training framework directly optimizes reasoning quality through structured reward signals, leading to substantial gains in both interpretability and localization accuracy. Empirical results show that GLOBE, using only 33K data, achieves performance comparable to or better than state-of-the-art methods trained on millions of samples.
Acknowledgements
The authors would like to thank the anonymous reviewers for their valuable comments. This work is funded by National Natural Science Foundation of China Grant (72371217, 62402414), the Guangzhou Industrial Informatics and Intelligence Key Laboratory No. 2024A03J0628, the Nansha Key Area Science and Technology Project No. 2023ZD003, and Project No. 2021JC02X191, and the Guangdong Basic and Applied Basic Research Foundation No. 2025A1515011994.
@@ A.1.1 Prompt Template
To ensure fairness and consistency in both training and evaluation, we standardize the input format for all LVLMs. Specifically, images are fed into the model along with a unified prompt template. This prompt is designed to guide the model toward reasoning about geographic cues in a consistent manner across different stages and model variants. During both fine-tuning and testing, we use the same prompt format to minimize prompt-induced variance and allow for fair comparison of model performance.
Prompt Template You are a geolocation expert. You are participating in a geolocation challenge. Based on the provided image: 1. Carefully analyze the image for clues about its location (architecture, signage, vegetation, terrain, etc.) 2. Think step-by-step about what country, and city this is likely to be in and why Your final answer includes these two lines somewhere in your response: country: [country name] city: [city name] You MUST output the thinking process in <think> </think> and give the answer in <answer> </answer> tags.
@@ A.1.3 Hyper-parameter Settings
A.2 Experimental Results
A.2.1 The performance of GLOBE across different conditions
Table A.2.1 presents geo-localization performance across three scene types (indoor, nature, and urban). We compare GLOBE-7B with Qwen2.5-VL-7B [24] and GeoReasoner-7B [28]. Across all conditions, GLOBE delivers consistently higher accuracy at every geographical level, demonstrating robust performance in diverse visual environments.
Table 7: Geo-localization performance comparison across different conditions.
Method MP16-Reason-Test (% @ km)
Street City Region Country Continent
1km 25km 200km 750km 2500km
[…]
Qwen2.5-VL-7B [24] 12.50 46.95 55.30 69.99 81.49
GeoReasoner-7B [28] 12.57 35.93 48.50 65.87 79.04
GLOBE-7B (Ours) 17.65 57.35 64.71 80.88 91.18
[…]
Qwen2.5-VL-7B [24] 8.61 42.77 60.07 72.62 80.68
GeoReasoner-7B [28] 5.10 35.71 48.98 67.35 78.57
GLOBE-7B (Ours) 13.95 55.81 81.40 90.70 97.67
[…]
Qwen2.5-VL-7B [24] 16.95 55.32 65.00 76.63 84.29
GeoReasoner-7B [28] 10.18 42.23 51.86 68.78 80.19
GLOBE-7B (Ours) 18.61 64.98 74.76 87.38 92.11
Best results are in blue.
Table 8: Geo-localization performance comparison on OSV-5M [83] (mini-3K).
Method OSV-5M [83] (mini-3K) (% @ km)
Street City Region Country Continent
1km 25km 200km 750km 2500km
ISNs [9] 0.00 1.07 6.77 22.04 44.01
GeoCLIP [15] 0.07 1.57 13.87 44.51 73.26
Qwen2.5-VL-7B [24] 0.00 0.87 5.14 19.81 40.55
InternVL3-8B [33] 0.00 0.73 5.27 19.81 44.01
GLOBE-7B (Ours) 0.00 1.87 14.04 45.01 74.16
Best results are in blue.
A.2.2 The performance of GLOBE on street-view images
Table 8 reports geo-localization performance on OSV-5M [83] (mini-3K), a benchmark consisting exclusively of street-view imagery. We compare GLOBE-7B against ISNs [9], GeoCLIP [15], and backbone-matched LVLMs such as Qwen2.5-VL-7B [24] and InternVL3-8B [33]. GLOBE achieves the best results at the city, region, country, and continent levels, demonstrating strong generalization to a domain outside its training distribution.
A.2.3 Failure cases
We categorize failure cases into two types: Error Reasoning and Right Reasoning, with representative examples illustrated in Table 9. Error Reasoning refers to cases where the model generates incorrect or irrelevant reasoning steps that misinterpret visual cues or contextual evidence. In contrast, Right Reasoning describes cases where the reasoning process is logically sound and visually grounded, yet the final prediction is incorrect (often due to dataset bias or overrepresentation of certain locations). Our analysis reveals two common patterns: ① visually similar features (e.g., domes, arches) leading to incorrect landmark attribution, and ② correct reasoning that is nevertheless biased toward locations more frequently represented in the training data.
Table 9: Representative failure cases.
Type IMG_ID Ground Truth Reasoning Prediction
eb_80_511397613.jpg Etterbeek, Belgium Terrain and Urban Setting: The grand scale of the architecture with large columns and arches resembles the Louvre Pyramid or the Arc de Triomphe area. Paris, France
98_80_2913796353.jpg Şehitkamil, Turkey Architecture: The minaret has a cylindrical structure with multiple levels featuring ornamental details like intricate carvings, wooden balconies with latticework, and stone masonry. The domed roof with a pointed tip and the presence of a crescent moon finial suggests Ottoman or Islamic architectural style. Baku, Azerbaijan
A.2.4 Inference efficiency
Table 10 reports the latency, throughput, and streaming metrics, with all models tested under the same hardware (single H20 GPU) and software settings (e.g., PyTorch, vLLM) as GLOBE. While GLOBE delivers stronger accuracy, it incurs higher average latency and lower throughput, reflecting the additional reasoning time required by larger LVLMs.
Table 10: Inference efficiency comparison between different baseline methods.
Model Concurrency Streaming Avg. Latency (s) Throughput (QPS) TTFT (ms) TPOT (ms)
GeoCLIP [15] 1 - 0.1364 7.3313 - -
RFM-YFCC [49] 1 - 0.5852 1.7088 - -
Qwen2.5-VL-7B [24] 1 No 2.9368 0.3405 - -
InternVL3-8B [33] 1 No 3.4886 0.2866 - -
GLOBE (InternVL3-8B [33]) 1 No 4.8742 0.2051 - -
GLOBE (InternVL3-8B [33]) 1 Yes 4.8597 0.2057 64.23 11.70
GLOBE (Qwen2.5-VL-7B [24]) 1 No 4.5684 0.2188 - -
GLOBE (Qwen2.5-VL-7B [24]) 1 Yes 4.5628 0.2191 34.67 11.72
GLOBE (Qwen2.5-VL-7B [24]) 8 No 5.1415 1.5045 - -
GLOBE (Qwen2.5-VL-7B [24]) 8 Yes 5.1589 1.4990 82.16 13.67
GLOBE (Qwen2.5-VL-7B [24]) 32 No 6.2479 4.5370 - -
GLOBE (Qwen2.5-VL-7B [24]) 32 Yes 6.2923 4.5602 161.40 17.12
A.2.5 Qualitative results
Figure 5 shows that GLOBE produces reasoning trajectories with improved coherence and interpretability. In particular, the model engages in structured reasoning to derive geo-location predictions, systematically incorporating diverse geographic cues such as architectural style, signage, vegetation, and other contextually informative elements.
Figure 5: Reasoning comparison of four different models (GPT-4.1 [86], GLOBE, Qwen2.5-VL-7B [24] with SFT, and InternVL3-78B [33]) on the same input image. Reliable visual cues identified by the models are marked in text.
A.2.6 Hyperparameters $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ in GRPO
The weights $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are fixed during training and were initially set to 0.2, 0.5, and 1, respectively. Since the primary objective is accurate prediction of city- and country-level locations, $\lambda_{3}$ (which directly supervises geo-localization accuracy) was assigned the highest weight. To mitigate hallucinations in the reasoning trajectory, $\lambda_{2}$ (consistency) was given a relatively high weight. In contrast, $\lambda_{1}$ (localizability), a binary score that evaluates whether the reasoning is geographically grounded, was assigned a smaller value to act as auxiliary regularization. As shown in Table 11, different weight combinations were tested, and the proposed setting (0.2, 0.5, 1) yielded the best performance.
Table 11: Performance comparison of weight selection ($\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$) for the GRPO framework.
$\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ MP16-Reason-Test (% @ km)
Street City Region Country Continent
1km 25km 200km 750km 2500km
1.0 0.5 0.2 17.63 59.96 72.11 84.87 91.55
1.0 1.0 1.0 17.67 59.94 71.83 84.80 91.20
0.2 0.5 1.0 17.99 62.85 73.83 86.68 92.52
The chosen configurations of $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are marked in bold.
A.2.7 Analysis of Reward Trajectories
The training dynamics of the three reward signals show generally upward trends (see Figure 6). The Localizability Reward steadily increases and gradually plateaus, showing consistent improvement in identifying informative inputs. Thanks to the pre-filtering of training data, both the Localizability Reward and the Geo-localization Accuracy Reward start with relatively high values and maintain strong performance even in the early training steps. The Visual Grounding Consistency Reward rises quickly during the initial stage before stabilizing, indicating that the model rapidly learns to associate visual entities with their corresponding locations. Overall, the trends of all three rewards suggest stable and effective learning throughout training.
Figure 6: Training dynamics of three rewards over global steps. From left to right: (a) Localizability Reward, (b) Visual Grounding Consistency Reward, and (c) Geo-localization Accuracy Reward. Each curve shows how the corresponding reward evolves as training progresses.
A.3 Limitation